if Rs750 amounts to rs 885 in three years at simple interest, what will be rs.1200 amount to in 3 1/2 years? (the rate is the same in both cases)
Answer by mananth(16946)

if Rs750 amounts to rs 885
885-750= 135 is the interest
let r be the rate of interest
135=(750*3*r)/100
(100*135)/(750*3)=r
r=6%
--------------
P=1200
r=6%
n=7/2 years (3 1/2)
I=pnr/100
I=(1200*7(/2)*6)/100
I=252
A= P+I
A=1200+252=Rs.1452
m.ananth@hotmail.ca